OpenStudy (anonymous):

Fractions: 4 1/4+2 1/8=

OpenStudy (anonymous):

51/8

OpenStudy (anonymous):

norbeybarajas: can u explain to me a lil how did u get that answer?

OpenStudy (anonymous):

multiply the whole fraction by the denominator of the fraction and you add the numerator

OpenStudy (anonymous):

4*4 =16+1 =17/4 and do the same with the other mixed number 2*8=16+1=17/8 17/4+17/8 =51/8

OpenStudy (anonymous):

can you give me another example please

OpenStudy (anonymous):

example: \[2 (4 / 3) + 5 (6/4 ) = 3*2+4=10/3 ; 5*4+6=26/4 ; 10/3 +26/4 = 59/6\]

OpenStudy (anonymous):

Leaving the same denominator

OpenStudy (anonymous):

thank you hope u can help me out with other math problems in the feature...

OpenStudy (anonymous):

ofcourse!!!